A body is projected up with a velocity equal to 3/4 of the escape velocity from the surface of the Earth. The height it reaches is (Radius of Earth is R) 

Grade:

2 Answers

Aman Bansal
12 years ago

Dear Siva,

vi=(3/4) of escape velocity from the surface of the earth = (3/4)*11200 = 8400 m/s
vf=0 m/s
t=time
g=9.8 m/s2
d=distance

vf=vi-gt
0=vi-gt
t=vi/g

now, d=(vf-vi)t/2
d=(vi)^2/2g

so,
d= 3600000 m = 3600km

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


Aman Bansal
12 years ago

Dear Siva,

Dont post ame query again and again.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert